OpenStudy (anonymous):

What is the mean absolute deviation for the data shown in the stem-and-leaf-plot? A. 6 B. 7 C. 20 D. 42

OpenStudy (igreen):

Where's the stem and leaf plot?

OpenStudy (anonymous):

there

OpenStudy (jacobciezki):

DONT HELP HIM

OpenStudy (igreen):

. . .

OpenStudy (anonymous):

Why not?

OpenStudy (jacobciezki):

DONT

OpenStudy (anonymous):

@jacobciezki get out

OpenStudy (anonymous):

o_o

OpenStudy (igreen):

This is our data: 9, 12, 18, 20, 26, 35 Find the mean. Add up the numbers and divide by how many numbers there are.

OpenStudy (jacobciezki):

DONT HELP HIM

OpenStudy (igreen):

I'm guessing you two are siblings or something..since both your names include "ciezki".

OpenStudy (anonymous):

yes

OpenStudy (jacobciezki):

GETOF NATHAN

OpenStudy (igreen):

If you don't give me a valid reason, I see no point in "not helping" him.

OpenStudy (anonymous):

Why don't you want him to be helped? @jacobciezki

OpenStudy (anonymous):

I agree.

OpenStudy (jacobciezki):

HE IS A BAD BOY

OpenStudy (igreen):

This is our data: 9, 12, 18, 20, 26, 35 Find the mean. Add up the numbers and divide by how many numbers there are. \(\sf\dfrac{9+12+18+20+26+35}{6}\) Simplify that

OpenStudy (anonymous):

i think its a or c

OpenStudy (igreen):

Can you simplify that up there?

OpenStudy (jacobciezki):

I THINK ITS DDD

OpenStudy (igreen):

@jacobciezki Please sort out your differences somewhere else.

OpenStudy (igreen):

Can you add 9 + 12 + 18 + 20 + 26 + 35? @nciezki.12132

OpenStudy (anonymous):

yes i got 120 and divided by 6 and got 20

OpenStudy (anonymous):

so its 20

OpenStudy (anonymous):

cc cc

OpenStudy (igreen):

Not yet @nciezki.12132

OpenStudy (anonymous):

ok

OpenStudy (igreen):

That's just the mean. Now we find the difference of the numbers from the mean. 9, 12, 18, 20, 26, 35 20 - 9 = ? 20 - 12 = ? 20 - 18 = ? 20 - 20 = ? 26 - 20 = ? 35 - 20 = ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill report back

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

how do you pull up pad

OpenStudy (igreen):

?

OpenStudy (anonymous):

the sheet

OpenStudy (igreen):

I don't know what you mean.

OpenStudy (anonymous):

42

OpenStudy (anonymous):

its 42

OpenStudy (igreen):

20 - 9 = 11 20 - 12 = 8 20 - 18 = 2 20 - 20 = 0 26 - 20 = 6 35 - 20 = 15 Add them up and divide by 6. \(\sf \dfrac{11 + 8 + 2 + 0 + 6 + 15}{6}\)

OpenStudy (igreen):

Oh, you're right..42..now divide by 6.

OpenStudy (anonymous):

7

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

That's our answer.